# Asymptotic behaviour of the Hitchin metric on the moduli space of Higgs bundles

### v2

@@ 1.3.5 Harmonic representatives of vertical infinitesimal deformations
The multiplication of $\tau$ on $L$ induces a $C^{\infty}$-section $F_{\tau}$ of End($\mathop{\rm End}\nolimits(E)\otimes\Omega^{0,1}_{X^{\circ}}$. It is helpful to prove that there exists $B_{1},B_{2}>0$ such that the following holds on $X\setminus\bigcup_{P\in D(\theta)}X_{P}(1/2)$:
@@
In general, for any complex manifold $Y$, let $Y^{\dagger}$ denote the conjugate of $Y$. There exist the natural isomorphism $T^{\ast}(X^{\dagger})\simeq(T^{\ast}X)^{\dagger}$.
@@
Let $P\in D(\theta)$. There exists an anti-holomorphic function $\beta_{P}$ on $\Sigma_{\theta|X(P)}$ such that ¯$\overline{\partial}\beta_{P}=\tau_{|\Sigma_{\theta|X_{P}}}$ and $\beta_{P}(Q)=0$ for any $Q\in\pi^{-1}(P)$. We may regard $\beta_{P}$ as a holomorphic function on $(\Sigma^{\dagger}_{t\theta})_{|X_{P}^{\dagger}}$. The multiplication of $\beta_{P}$ on $L^{\dagger}_{t}$ induces the endomorphism $F_{(E^{\dagger}_{t},t\theta^{\dagger}_{h_{t}}),\beta_{P}}$ of $E^{\dagger}_{t}$. It satisfies
@@
By the asymptotic orthogonality in [20], we obtain ${\tt V}_{P}(\tau,t)-F_{\tau}=O\bigl{(}e^{-\beta t}\bigr{)}$ on $X_{P}(1)\setminus X_{P}(1/2)$ for some $\beta>0$.
@@
We have $F_{(E^{\dagger}_{t},t\theta^{\dagger}_{h_{t}}),\beta_{P}}=(F_{\overline{\beta}_{P}})^{\dagger}_{h_{t}}$, where $F_{\overline{\beta}_{P}}$ is the holomorphic endomorphism of $(E,\theta)_{|X_{P}}$ induced by the holomorphic function $\beta_{P}$ on $\Sigma_{\theta|X_{P}}$.
@@
$\tau^{\prime}=\tau-\sum_{P\in D(\theta)}d\bigl{(}\chi_{P}\beta_{P}\bigr{)}.$
We obtain the $C^{\infty}$-section $F_{\tau^{\prime}}$ of $\mathop{\rm End}\nolimits(E)\otimes\Omega^{0,1}$ such that $[F_{\tau^{\prime}}]=\iota^{{\tt v}}_{t}(\tau)$. We set
@@
Then, ${\tt V}^{\prime}(\tau,t)$ satisfies (7) for some $B_{i}>0$. We have (¯$(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)^{\ast}_{h_{t},g_{X}}{\tt V}^{\prime}(\tau,t)=O\bigl{(}e^{-\epsilon t}\bigr{)}\|\tau\|$ on $X$ for some $\epsilon>0$. Moreover, $\mathop{\rm tr}\nolimits\bigl{(}(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)^{\ast}_{h_{t},g_{X}}{\tt V}^{\prime}(\tau,t)\bigr{)}=0$ holds. Then, for each $t>0$, there uniquely exists $\gamma(\tau,t)\in A^{0}(\mathop{\rm End}\nolimits(E))$ such that $\mathop{\rm tr}\nolimits\gamma(\tau,t)=0$ and
(¯$(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)^{\ast}_{h_{t},g_{X}}\circ(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)\gamma(\tau,t)=(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)^{\ast}_{h_{t},g_{X}}{\tt V}^{\prime}(\tau,t).$
Again, by using the estimate for the first eigenvalues of the Laplacian operators (¯$(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)^{\ast}_{h_{t},g_{X}}\circ(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)$ (see Proposition 2.42) and the regularity for solutions of the Poisson equation (see Proposition 2.36), we obtain $|\gamma(\tau,t)|_{h_{t}}+|(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)\gamma(\tau,t)|_{h_{t},g_{X}}\leq B_{1}\exp(-B_{2}t)\|\tau\|$ for some $B_{i}>0$. Because ${\tt V}(\tau,t)={\tt V}^{\prime}(\tau,t)-(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)\gamma(\tau,t)$, we obtain the estimate (7) for ${\tt V}(\tau,t)$.
@@ 1.3.6 Estimate of pairings with vertical infinitesimal deformations
Let us explain an outline of the proof of the estimate (10). Because (¯$(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)^{\ast}_{h_{t},g_{X}}{\tt V}(\tau,t)=0$, we obtain
$\sqrt{-1}\int_{X}\Bigl{(}h_{t}\bigl{(}{\tt V}(\tau,t)^{1,0},{\tt V}(\tau,t)^{1,0}\bigr{)}-h_{t}\bigl{(}{\tt V}(\tau,t)^{0,1},{\tt V}(\tau,t)^{0,1}\bigr{)}\Bigr{)}=-\sqrt{-1}\int_{X}\Bigl{(}h_{t}\bigl{(}F_{\tau^{\prime}},{\tt V}(\tau,t)^{0,1}\bigr{)}\Bigr{)}\\ =-\sqrt{-1}\int_{X}\mathop{\rm tr}\nolimits\Bigl{(}F_{\tau^{\prime}}\cdot\bigl{(}{\tt V}(\tau,t)^{0,1}\bigr{)}^{\dagger}_{h_{t}}\Bigr{)}.$ (12)
@@
$\sqrt{-1}\int_{\Sigma_{\theta}}\tau\overline{\tau}=\sqrt{-1}\int_{\Sigma_{\theta}}\tau^{\prime}\overline{\tau}=\sqrt{-1}\int_{X}\mathop{\rm tr}\nolimits\bigl{(}F_{\tau^{\prime}}F_{\overline{\tau}}\bigr{)}.$
By the estimate (7) and the asymptotic orthogonality in [20], the following holds on $X\setminus\bigcup X_{P}(1/2)$ for some $B_{3},B_{4}>0$:
@@ 1.3.6.1 Acknowledgements
This study grew out of my efforts to understand the attractive works [4, 5, 19]. I hope that this would be useful in the further study of this subject. I thank Qiongling Li and Szilard Szabo for discussions and collaborations. This study is partially based on the joint works [17, 23]. I am grateful to Gao Chen, Laura Fredrickson, Nianzi Li, Rafe Mazzeo and Siqi He for some discussions. In particular, I appreciate Fredrickson and Mazzeo for sincerely replying to my inquiries regarding Remark 1.2. I am grateful to Fredrickson, Mazzeo and Szabo for helpful comments to this manuscript. I thank Hitoshi Fujioka for discussions. I thank Yoshifumi Tsuchimoto and Akira Ishii for their constant encouragements.
@@
I am partially supported by the Grant-in-Aid for Scientific Research (A) (No. 21H04429), the Grant-in-Aid for Scientific Research (A) (No. 22H00094), the Grant-in-Aid for Scientific Research (A) (No. 23H00083), and the Grant-in-Aid for Scientific Research (C) (No. 20K03609), Japan Society for the Promotion of Science. I am also partially supported by the Research Institute for Mathematical Sciences, an International Joint Usage/Research Center located in Kyoto University.
@@ 2.1 Some notation
Let $(E,\theta)$ be a Higgs bundle on $X$, i.e., $E$ denotes a holomorphic vector bundle on $X$, and $\theta$ denotes a holomorphic section of $\mathop{\rm End}\nolimits(E)\otimes K_{X}$. The spectral curve of the Higgs bundle $(E,\theta)$ is denoted by $\Sigma_{\theta}$, i.e., $\Sigma_{\theta}$ denotes the 00-scheme of the section $\det(\eta\mathop{\rm id}\nolimits_{E}-\theta)$ of $\pi_{0}^{\ast}\bigl{(}K_{X}^{\otimes\mathop{\rm rank}\nolimits(E)}\bigr{)}$. For any subset $U\subset X$, we set $\Sigma_{\theta|U}=\Sigma_{\theta}\times_{X}U$. We set
@@ 2.1.2 Filtrations
There exists a decreasing filtration $\mathcal{F}^{j}(E_{Q|P})$ $(j=0,\ldots,r_{Q}-1)$ obtained as the images of
@@ 2.2.1 Deformation complexes of Higgs bundles
Here, the first term sits in the degree 00.
@@ 2.2.2 The morphism $\mathop{\rm ad}\nolimits(\theta)$ in the case where the spectral curve is smooth
Proof   Let us prove (15). Let $P\in X$ with a holomorphic coordinate neighbourhood $(X_{P},z_{P})$ as in §2.1.1. We obtain the coordinate $(z,\xi)$ on $T^{\ast}X_{P}$. There exists the decomposition (14) of $(E,\theta)_{|X_{P}}$. Each connected component $(\Sigma_{\theta})_{Q}$ is smooth. The eigenvalues of $f_{Q|P}$ are $\xi(Q)$. Hence, it is enough to consider the case where $|\Sigma_{\theta|P}|=1$ and $X=X_{P}$. Moreover, we may assume that $\Sigma_{\theta|P}=\{0\}$ in $T_{P}^{\ast}X$ by replacing $\theta$ with $\theta-\frac{1}{r}\mathop{\rm tr}\nolimits(\theta)\mathop{\rm id}\nolimits_{E}$. We set $\widetilde{X}=\Sigma_{\theta}$. Let $\pi:\widetilde{X}\to X$ denote the projection. We identify $T^{\ast}X=X\times{\mathbb{C}}$ by using the coordinate $z_{P}$. It induces a holomorphic coordinate of $\widetilde{X}$. There exists an isomorphism $E\simeq\pi_{\ast}(\mathcal{O}_{\widetilde{X}})$ under which the endomorphism $f$ is induced by the multiplication of $\xi$ on $\mathcal{O}_{\widetilde{X}}$. We obtain the frame $v_{j}=\pi_{\ast}(\xi^{j})$ $(j=0,\ldots,r-1)$ of $E$. Because $f_{P}(v_{j|P})=v_{j+1|P}$ $(j<r-1)$ and $f_{P}(v_{r-1|P})=0$, we obtain (15). The other claims follow from (15).
@@
Let $N_{Q}$ denote the nilpotent part of $f_{Q|P}$. Then, we have $N_{Q}(\mathcal{F}^{j}(E_{P|Q}))\subset\mathcal{F}^{j+1}(E_{P|Q})$ for $j=0,\ldots,r_{Q}-1$. Moreover, $N_{Q}$ induces an isomorphism $\mathop{\rm Gr}\nolimits^{\mathcal{F}}_{j}(E_{Q|P})\simeq\mathop{\rm Gr}\nolimits^{\mathcal{F}}_{j+1}(E_{Q|P})$ for $j=0,\ldots,r_{Q}-1$.
@@ 2.2.3 Deformation complex in the case where $\Sigma_{\theta}$ is smooth
Let $P\in X^{\circ}$. There exist a coordinate neighbourhood $(X_{P},z_{P})$ and a decomposition of the Higgs bundle (14) as in §2.1.1. Let $\psi$ be any holomorphic section of $\mathop{\rm End}\nolimits(E)\otimes K_{X}$ on $X_{P}$. We have the decomposition $\psi=\sum\psi_{Q,Q^{\prime}}$, where $\psi_{Q,Q^{\prime}}\in\mathop{\rm Hom}\nolimits(E_{Q^{\prime}},E_{Q})\otimes K_{X}$. We may naturally regard $\sum_{Q}\psi_{Q,Q}$ as a section of $\pi_{\ast}(K_{\Sigma_{\theta}})_{|X_{P}}$. This induces an isomorphism $\rho_{P}:\mathop{\rm Cok}\nolimits(\mathop{\rm ad}\nolimits\theta)_{|X_{P}}\simeq\pi_{\ast}(K_{\Sigma_{\theta}})_{|X_{P}}$. By considering the isomorphisms $\rho_{P}$ for any $P\in X^{\circ}$, we obtain the isomorphism $\rho_{X^{\circ}}:\mathop{\rm Cok}\nolimits\bigl{(}\mathop{\rm ad}\nolimits\theta\bigr{)}_{|X^{\circ}}\simeq\pi_{\ast}(K_{\Sigma_{\theta}})_{|X^{\circ}}$. Let us prove that $\rho_{X^{\circ}}$ extends to an isomorphism $\rho_{X}:\mathop{\rm Cok}\nolimits(\mathop{\rm ad}\nolimits\theta)\simeq\pi_{\ast}(K_{\Sigma_{\theta}})$. It is enough to consider the case where $X$ is a neighbourhood $U$ of 00 in ${\mathbb{C}}$ and $\Sigma_{\theta|0}=\{0\}$. Let $f$ be the endomorphism of $E$ determined by $\theta=f\,dz$. We set $U^{\circ}=U\setminus\{0\}$ and $\widetilde{U}=\Sigma_{\theta}$.
@@
The isomorphism $\rho_{U^{\circ}}$ extends to a morphism $\rho_{U}:\mathop{\rm Cok}\nolimits(\mathop{\rm ad}\nolimits\theta)\to\pi_{\ast}(K_{\widetilde{U}})$.
@@
Let ${\boldsymbol{u}}=(u_{0},\ldots,u_{r-1})$ and ${\boldsymbol{v}}=(v_{0},\ldots,v_{r-1})$. For any holomorphic section $g$ of $\mathop{\rm End}\nolimits(E)\otimes K_{U}$ on $U$, we obtain the matrix valued $1$-form $G$ on $U$ determined by $g({\boldsymbol{u}})={\boldsymbol{u}}G$. Let $B$ be the matrix valued holomorphic function on $U^{(r)}$ determined by ${\boldsymbol{v}}=\varphi_{r}^{\ast}({\boldsymbol{u}})B$. We have
@@
Proof   Let $\mathfrak{s}\mathfrak{l}(2,{\mathbb{C}})$ denote the space of $(2\times 2)$-matrices with 00. Let
@@
There exists a representation $\kappa$ of $\mathfrak{s}\mathfrak{l}(2,{\mathbb{C}})$ on $V$ such that $\kappa(f)=N$. There exists the decomposition $\mathop{\rm End}\nolimits(V)={\mathbb{C}}\cdot\mathop{\rm id}\nolimits_{V}\oplus\mathop{\rm End}\nolimits_{0}(V)$, where $\mathop{\rm End}\nolimits_{0}(V)$ denotes the space of the trace-free endomorphisms of $V$. It is compatible with the induced $\mathfrak{s}\mathfrak{l}(2,{\mathbb{C}})$-action. It is well known that $\mathop{\rm End}\nolimits_{0}(V)$ does not contain a trivial representation of $\mathfrak{s}\mathfrak{l}(2,{\mathbb{C}})$, and $N^{j}$ $(j=1,\ldots,r-1)$ are contained in $\mathop{\rm End}\nolimits_{0}(V)$. Then, the claim is clear.
@@
$\pi_{\ast}(K_{\widetilde{U}})$ equals the image of $\rho_{U}$.
Proof   Let $\nu$ be a holomorphic $1$-form on $\widetilde{U}$. The multiplication of $\nu$ induces a section $F_{\nu}$ of $\mathop{\rm End}\nolimits(E)\otimes K_{U}(\{0\})$. We have $[f,F_{\nu}]=0$. It implies $[f_{|0},\mathop{\rm Res}\nolimits_{0}(F_{\nu})]=0$. It is easy to see that $\mathop{\rm Res}\nolimits_{0}(F_{\nu})\in\sum_{j=1}^{r-1}{\mathbb{C}}f_{|0}^{j}$. By Lemma 2.6, there exists $g_{0}\in\mathop{\rm End}\nolimits(E_{|0})$ such that $[f,g_{0}]=\mathop{\rm Res}\nolimits_{0}(F_{\nu})$. Let $g$ be a holomorphic section of $\mathop{\rm End}\nolimits(E)$ such that $g_{|0}=g_{0}$. Then, $\widetilde{F}=F_{\nu}-[\theta,z^{-1}g]$ is a holomorphic section of $\mathop{\rm End}\nolimits(E)\otimes K_{U}$, and $\rho_{U}(\widetilde{F})=\pi_{\ast}(\nu)$. Thus, we obtain Lemma 2.7 and Proposition 2.4.

### v3

@@ 1.3.5 Harmonic representatives of vertical infinitesimal deformations
The multiplication of $\tau$ on $L$ induces a $C^{\infty}$-section $F_{\tau}$ of $\mathop{\rm End}\nolimits(E)\otimes\Omega^{0,1}_{X^{\circ}}$. It is helpful to prove that there exist $B_{1},B_{2}>0$ such that the following holds on $X\setminus\bigcup_{P\in D(\theta)}X_{P}(1/2)$:
@@
In general, for any complex manifold $Y$, let $Y^{\dagger}$ denote the conjugate of $Y$. There exists the natural isomorphism $T^{\ast}(X^{\dagger})\simeq(T^{\ast}X)^{\dagger}$.
@@
Let $P\in D(\theta)$. There exists an anti-holomorphic function $\beta_{P}$ on $\Sigma_{\theta|X(P)}$ such that $\overline{\partial}\beta_{P}=\tau_{|\Sigma_{\theta|X_{P}}}$ and $\beta_{P}(Q)=0$ for any $Q\in\pi^{-1}(P)$. We may regard $\beta_{P}$ as a holomorphic function on $(\Sigma^{\dagger}_{t\theta})_{|X_{P}^{\dagger}}$. The multiplication of $\beta_{P}$ on $L^{\dagger}_{t}$ induces the endomorphism $F_{(E^{\dagger}_{t},t\theta^{\dagger}_{h_{t}}),\beta_{P}}$ of $E^{\dagger}_{t}$. It satisfies
@@
By the asymptotic orthogonality in [20], we obtain ${\tt V}_{P}(\tau,t)-F_{\tau}=O\bigl(e^{-\epsilon t}\bigr)$ on $X_{P}(1)\setminus X_{P}(1/2)$ for some $\epsilon>0$.
@@
We have $F_{(E^{\dagger}_{t},t\theta^{\dagger}_{h_{t}}),\beta_{P}}=(F_{\overline{\beta}_{P}})^{\dagger}_{h_{t}}$, where $F_{\overline{\beta}_{P}}$ is the holomorphic endomorphism of $(E,\theta)_{|X_{P}}$ induced by the holomorphic function $\overline{\beta}_{P}$ on $\Sigma_{\theta|X_{P}}$.
@@
$\tau^{\prime}=\tau-\sum_{P\in D(\theta)}\overline{\partial}\bigl(\chi_{P}\beta_{P}\bigr).$
We obtain the $C^{\infty}$-section $F_{\tau^{\prime}}$ of $\mathop{\rm End}\nolimits(E)\otimes\Omega^{0,1}$ such that $F_{\tau^{\prime}}]=\iota^{{\tt v}}_{t}(\tau)$. We set
@@
Then, ${\tt V}^{\prime}(\tau,t)$ satisfies (7) for some $B_{i}>0$. We have $(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)^{\ast}_{h_{t},g_{X}}{\tt V}^{\prime}(\tau,t)=O\bigl(e^{-\epsilon t}\bigr)\|\tau\|$ on $X$ for some $\epsilon>0$. Moreover, $\mathop{\rm tr}\nolimits\bigl((\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)^{\ast}_{h_{t},g_{X}}{\tt V}^{\prime}(\tau,t)\bigr)=0$ holds. Then, for each $t>0$, there uniquely exists $\gamma(\tau,t)\in A^{0}(\mathop{\rm End}\nolimits(E))$ such that $\mathop{\rm tr}\nolimits\gamma(\tau,t)=0$ and
$(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)^{\ast}_{h_{t},g_{X}}\circ(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)\gamma(\tau,t)=(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)^{\ast}_{h_{t},g_{X}}{\tt V}^{\prime}(\tau,t).$
Again, by using the estimate for the first eigenvalues of the Laplacian operators $(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)^{\ast}_{h_{t},g_{X}}\circ(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)$ (see Proposition 2.41) and the regularity for solutions of the Poisson equation (see Proposition 2.35), we obtain $|\gamma(\tau,t)|_{h_{t}}+|(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)\gamma(\tau,t)|_{h_{t},g_{X}}\leq B_{1}\exp(-B_{2}t)\|\tau\|$ for some $B_{i}>0$. Because ${\tt V}(\tau,t)={\tt V}^{\prime}(\tau,t)-(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)\gamma(\tau,t)$, we obtain the estimate (7) for ${\tt V}(\tau,t)$.
@@ 1.3.6 Estimate of pairings with vertical infinitesimal deformations
Let us explain an outline of the proof of the estimate (10). Because $(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)^{\ast}_{h_{t},g_{X}}{\tt V}(\tau,t)=0$, we obtain
$\sqrt{-1}\int_{X}\Bigl(h_{t}\bigl({\tt V}(\tau,t)^{1,0},{\tt V}(\tau,t)^{1,0}\bigr)-h_{t}\bigl({\tt V}(\tau,t)^{0,1},{\tt V}(\tau,t)^{0,1}\bigr)\Bigr)=-\sqrt{-1}\int_{X}\Bigl(h_{t}\bigl(F_{\tau^{\prime}},{\tt V}(\tau,t)^{0,1}\bigr)\Bigr)\\ =-\sqrt{-1}\int_{X}\mathop{\rm tr}\nolimits\Bigl(F_{\tau^{\prime}}\cdot\bigl({\tt V}(\tau,t)^{0,1}\bigr)^{\dagger}_{h_{t}}\Bigr)=-\sqrt{-1}\int_{X\setminus\bigcup X_{P}(1/2)}\mathop{\rm tr}\nolimits\Bigl(F_{\tau^{\prime}}\cdot\bigl({\tt V}(\tau,t)^{0,1}\bigr)^{\dagger}_{h_{t}}\Bigr)$ (12)
@@
$\sqrt{-1}\int_{\Sigma_{\theta}}\tau\overline{\tau}=\sqrt{-1}\int_{\Sigma_{\theta}}\tau^{\prime}\overline{\tau}=\sqrt{-1}\int_{X}\mathop{\rm tr}\nolimits\bigl(F_{\tau^{\prime}}F_{\overline{\tau}}\bigr)=\sqrt{-1}\int_{X\setminus\bigcup X_{P}(1/2)}\mathop{\rm tr}\nolimits\bigl(F_{\tau^{\prime}}F_{\overline{\tau}}\bigr).$
Note that the support of $F_{\tau^{\prime}}$ is contained in $X\setminus\bigcup X_{P}(1/2)$. By the estimate (7) and the asymptotic orthogonality in [20], the following holds on $X\setminus\bigcup X_{P}(1/2)$ for some $B_{3},B_{4}>0$:
@@ 1.3.6.1 Acknowledgements
I thank the reviewer for his/her thorough and careful reading and helpful comments to improve this manuscript. This study grew out of my efforts to understand the attractive works [4, 5, 19]. I hope that this would be useful in the further study of this subject. I thank Qiongling Li and Szilard Szabo for discussions and collaborations. This study is partially based on the joint works [17, 24]. I am grateful to Gao Chen, Laura Fredrickson, Nianzi Li, Rafe Mazzeo and Siqi He for some discussions. In particular, I appreciate Fredrickson and Mazzeo for sincerely replying to my inquiries regarding Remark 1.2. I am grateful to Fredrickson, Mazzeo and Szabo for helpful comments to this manuscript. I thank Hitoshi Fujioka for discussions. I thank Yoshifumi Tsuchimoto and Akira Ishii for their constant encouragements.
@@
I am partially supported by the Grant-in-Aid for Scientific Research (A) (No. 21H04429), the Grant-in-Aid for Scientific Research (A) (No. 22H00094), the Grant-in-Aid for Scientific Research (A) (No. 23H00083), the Grant-in-Aid for Scientific Research (C) (No. 20K03609), the Grant-in-Aid for Scientific Research (C) (No. 25K06973), Japan Society for the Promotion of Science. I am also partially supported by the Research Institute for Mathematical Sciences, an International Joint Usage/Research Center located in Kyoto University.
@@ 2.1 Some notation
Let $(E,\theta)$ be a Higgs bundle on $X$, i.e., $E$ denotes a holomorphic vector bundle on $X$, and $\theta$ denotes a holomorphic section of $\mathop{\rm End}\nolimits(E)\otimes K_{X}$. The spectral curve of the Higgs bundle $(E,\theta)$ is denoted by $\Sigma_{\theta}$, i.e., $\Sigma_{\theta}$ denotes the 0-scheme of the section $\det(\eta\mathop{\rm id}\nolimits_{E}-\theta)$ of $\pi_{0}^{\ast}\bigl(K_{X}^{\otimes\mathop{\rm rank}\nolimits(E)}\bigr)$. For any subset $U\subset X$, we set $\Sigma_{\theta|U}=\Sigma_{\theta}\times_{X}U$. We set
@@ 2.1.2 Filtrations
When $\Sigma_{\theta}$ is smooth and there exists a line bundle $L$ on $\Sigma_{\theta}$ such that $E=\pi_{\ast}(L)$, there exists a decreasing filtration $\mathcal{F}^{j}(E_{Q|P})$ $(j=0,\ldots,r_{Q}-1)$ obtained as the images of
@@ 2.2.1 Deformation complexes of Higgs bundles
Here, the first term sits in the degree 0.
@@ 2.2.2 The morphism $\mathop{\rm ad}\nolimits(\theta)$ in the case where the spectral curve is smooth
[…]
@@
Let $\mathcal{F}$ be the filtration in §2.1.2. Let $N_{Q}$ denote the nilpotent part of $f_{Q|P}$. Then, we have $N_{Q}(\mathcal{F}^{j}(E_{P|Q}))\subset\mathcal{F}^{j+1}(E_{P|Q})$ for $j=0,\ldots,r_{Q}-1$. Moreover, $N_{Q}$ induces an isomorphism $\mathop{\rm Gr}\nolimits^{\mathcal{F}}_{j}(E_{Q|P})\simeq\mathop{\rm Gr}\nolimits^{\mathcal{F}}_{j+1}(E_{Q|P})$ for $j=0,\ldots,r_{Q}-1$.
@@ 2.2.3 Deformation complex in the case where $\Sigma_{\theta}$ is smooth
Let $P\in X^{\circ}$. There exist a coordinate neighbourhood $(X_{P},z_{P})$ and a decomposition of the Higgs bundle (14) as in §2.1.1. Let $\psi$ be any holomorphic section of $\mathop{\rm End}\nolimits(E)\otimes K_{X}$ on $X_{P}$. We have the decomposition $\psi=\sum\psi_{Q,Q^{\prime}}$, where $\psi_{Q,Q^{\prime}}\in\mathop{\rm Hom}\nolimits(E_{Q^{\prime}},E_{Q})\otimes K_{X}$. We may naturally regard $\sum_{Q}\psi_{Q,Q}$ as a section of $\pi_{\ast}(K_{\Sigma_{\theta}})_{|X_{P}}$. This induces an isomorphism $\rho_{P}:\mathop{\rm Cok}\nolimits(\mathop{\rm ad}\nolimits\theta)_{|X_{P}}\simeq\pi_{\ast}(K_{\Sigma_{\theta}})_{|X_{P}}$. By considering the isomorphisms $\rho_{P}$ for any $P\in X^{\circ}$, we obtain the isomorphism $\rho_{X^{\circ}}:\mathop{\rm Cok}\nolimits\bigl(\mathop{\rm ad}\nolimits\theta\bigr)_{|X^{\circ}}\simeq\pi_{\ast}(K_{\Sigma_{\theta}})_{|X^{\circ}}$. Let us prove that $\rho_{X^{\circ}}$ extends to an isomorphism $\rho_{X}:\mathop{\rm Cok}\nolimits(\mathop{\rm ad}\nolimits\theta)\simeq\pi_{\ast}(K_{\Sigma_{\theta}})$. It is enough to consider the case where $X$ is a neighbourhood $U$ of 0 in $\mathbb{C}$} and $\Sigma_{\theta|0}=\{0\}$. Let $f$ be the endomorphism of $E$ determined by $\theta=f\,dz$. We set $U^{\circ}=U\setminus\{0\}$ and $\widetilde{U}=\Sigma_{\theta}$.
@@
[…]
@@
Let ${\boldsymbol{u}}=(u_{0},\ldots,u_{r-1})$ and ${\boldsymbol{v}}=(v_{0},\ldots,v_{r-1})$. For any holomorphic section $g$ of $\mathop{\rm End}\nolimits(E)\otimes K_{U}$ on $U$, we obtain the matrix valued holomorphic $1$-form $G$ on $U$ determined by $g({\boldsymbol{u}})={\boldsymbol{u}}G$. Let $B$ be the matrix valued meromorphic function on $U^{(r)}$ determined by ${\boldsymbol{v}}=\varphi_{r}^{\ast}({\boldsymbol{u}})B$. We have
@@
Proof  Let $\mathfrak{s}\mathfrak{l}(2,{\mathbb{C}})$ denote the Lie algebra of trace-free $(2\times 2)$-matrices. Let
@@
There exists an irreducible representation $\kappa$ of $\mathfrak{s}\mathfrak{l}(2,{\mathbb{C}})$ on $V$ such that $\kappa(f)=N$. There exists the decomposition $\mathop{\rm End}\nolimits(V)={\mathbb{C}}\cdot\mathop{\rm id}\nolimits_{V}\oplus\mathop{\rm End}\nolimits_{0}(V)$, where $\mathop{\rm End}\nolimits_{0}(V)$ denotes the space of the trace-free endomorphisms of $V$. It is compatible with the induced $\mathfrak{s}\mathfrak{l}(2,{\mathbb{C}})$-action. It is well known that $\mathop{\rm End}\nolimits_{0}(V)$ does not contain a trivial representation of $\mathfrak{s}\mathfrak{l}(2,{\mathbb{C}})$, and $N^{j}$ $(j=1,\ldots,r-1)$ are contained in $\mathop{\rm End}\nolimits_{0}(V)\cap\mathop{\rm Ker}\nolimits\mathop{\rm ad}\nolimits(N)$. Then, the claim is clear.
@@
[…]
Proof  Let $\nu$ be a holomorphic $1$-form on $\widetilde{U}$. The multiplication of $\nu$ induces a section $F_{\nu}$ of $\mathop{\rm End}\nolimits(E)\otimes K_{U}(\{0\})$. We have $f,F_{\nu}]=0$. It implies $f_{|0},\mathop{\rm Res}\nolimits_{0}(F_{\nu})]=0$. It is easy to see that $\mathop{\rm Res}\nolimits_{0}(F_{\nu})\in\sum_{j=1}^{r-1}{\mathbb{C}}f_{|0}^{j}$. By Lemma 2.6, there exists $g_{0}\in\mathop{\rm End}\nolimits(E_{|0})$ such that $f,g_{0}]=\mathop{\rm Res}\nolimits_{0}(F_{\nu})$. Let $g$ be a holomorphic section of $\mathop{\rm End}\nolimits(E)$ such that $g_{|0}=g_{0}$. Then, $\widetilde{F}=F_{\nu}-[\theta,z^{-1}g]$ is a holomorphic section of $\mathop{\rm End}\nolimits(E)\otimes K_{U}$, and $\rho_{U}(\widetilde{F})=\pi_{\ast}(\nu)$. Thus, we obtain Lemma 2.7 and Proposition 2.4.